Tom was paid 4.50 an hour fot painting a fence and 6.50 an hour to mowing grass. He worked 7 hours and earned 39.50. How many hours did it take him to paint the fence?

Answer by ReadingBoosters(3246)

4.50/hr for fence
6.50/hr for grass
4.50x + 6.50y = 39.50
x+y=7
Substitute into the first equation
4.50(7-y) + 6.50y = 39.50
31.50 - 4.50y + 6.50y = 39.50
31.50 + 2y = 39.50
2y = 39.50 - 31.50
2y = 8
y = 4 hours mowing grass
x = 7-4= 3 hours painting the fence
Proof
4.5(3) + 6.5(4) =
13.50 + 26 = 39.50
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.